Log

Date and time Initials Description Remarks
2.9, 19:02 TB BGO HV to 445V - started run 1441213330 Monitoring shows a little gain over 410V
Best range may be ~460-470V
2.9, 15:40 TB Set BGO HV back to 410V - started run 1441212547 Re-checking results obtained in 1441190945
Amplifier at x100, BGO trigger threshold at 500mV
Monitoring looks as before - BGO has rapid turn on from 410V -> 480V
2.9, 15:40 TB Set BGO HV to 480V - started run 1441205369 Trigger threshold still 500mV
2.9, 15:40 TB Set BGO HV to 550V - started run 1441201195 Trigger threshold still 500mV
2.9, 12:45 TB Set BGO amp back to x100 and taken run 1441190945 BGO HV at 410V, trigger threshold increased to 500mV
2.9, 12:30 TB Set BGO amp from x100->x200 and taken run 1441189955 BGO trigger threshold increased to 130mV
2.9, 11:43 TB Set BGO amp from x50->x100 and taken run 1441189261  
2.9, 11:43 TB Set BGO amp from x5->x50 and taken run 1441187265 BGO trigger threshold increased to 50mV
2.9, 11:30 TB Taken run 1441185871 for BGO noise/cosmic data Triggering from BGO edge. BGO signal in ADC_0, attenuated copy in ADC_1
1.9, 18:00 TB, JP Connected up BGO Should take run of BGO pulses for cosmic/noise calibration
1.9, 18:00 TB, JP Covered LG3 to test light leak measured normal pedestal
1.9, 18:00 TB, JP Found LG4 connected to faulty patch panel connection (ch0) Switched to high patch panel - normal pedestal values read
1.9, 18:00 TB, JP Pedestal program for v785 pinned to 16 channels for v785N module Note channel bits offset by 1 to 'plain' v785
1.9, 18:00 TB, JP Covered LG3 to test light leak measured normal pedestal
29.8, 21:40 TB Taken runs 1440876475 - 1440877156 with pulser to calibrate DWC1 & 2
29.8, 19:50 TB, JP Taken runs 1440866379 for LG4 and 1440868586 for LG3 Blocks were moved behind veto SC1 in theoretical beam center
LG3 seems to have some activity over the pedestal charge - evident in the 1440866298.ped file
29.8, 15:30 TB, JP Taken run 1440856097 with veto infront of BGO calo
29.8, 15:20 TB, JP Adjusted beam focus to 4.5m(run 1440853570) - then 7m(run 1440854958) Calo off - no veto scintillator
29.8, 11:00 TB, JP Fixed DWC0 & 1 DWC0-R was disconnected from the fixed descriminator
DWC1-R was attenuated due to a break in the cable (4-08) - borrowed cable 4-10 for now.
28.8, 20:50 TB, JP Taken run 1440783291 Trigger SC0&!SC1&SC2&CH0
28.8, 19:16 TB, JP Taken runs 144078{1966,0173} Trigger SC0&SC2&CH0
28.8, 14:00 TB, JP Received magnet MNP17 Michael Lazzeroni supervised install, Lau Gattinon and Michael Jeckel visited installation later
28.8, 11:00 TB, JP Visited cable service to get 2 more 20m cables Should be complete monday
28.8, 10:00 TB, CD, JP Continued re-cabling detectors after relocation TODO: update cable map
27.9, 16:00 TB, CD, JP Received stands for magnet, re-located DESY table Relocated calorimeters onto DESY table
Started re-cabling towards final configuration
27.8, 14:00 TB Measuring detector rates with beam Rates per spill:
SC0 ~22k
SC1 ~20k
SC2 ~13k
CH0
CH1
26.8, 6:30 TB Put DWC2 into gas supply in serial with DWC1
Cabled DWC2 to 'P' loom
Cabled up new discriminators
Need to check timing on new scintillators
26.8, Afternoon TB,JP Switched faulty discriminator for 2 working ones. Debugged problem with timer pulser not triggering discriminator - used another pulser - works
25.8, Afternoon TB,JP,MJ Mounted DWC2 in beamline
24.8, 8:00 TB, JP, PA Taken run 1440436068 with no LGs but all BGO, SC, CH BGO amp at 5x, voltage at 410V
Only CH1 in trigger due to faulty discriminator
24.8, Morning MJ, PA Ran RS232 cable from control room to zone Later, tested communication to controller unit
21.8, 20:00 TB, JP Taken run 1440179961 with V785 (BGO was in the beam)
21.8, 17:59 TB, JP Taken run 1440172747 with SC0 + SC3 trigger Same conditions otherwise
21.8, 17:30 TB, JP, PA Set up trigger logic (CH0 + CH1) + SC0 + SC3
Recorded run 1440170076
SC0 signal (TDC0-08) seems missing - may be too early for TDC window
21.8, 10:52 TB, PA Installed BGO and support crate, veto scintillator
20.8, 19:07 MJ, CD Have seen the beam with the D05 TimePix Problems
If pixman creates a png file the file does not contain the raw data
Currently the png file is not created because pixman tries to write /afs/cern.png. We have to discuss this with the experts
20.8, 17:07 TB, JP, PA Recorded run 1440082502 25k events at 3GeV negative beam.
Cherenkov pressures at 2 atm. LG0 in beam center, LG4 disconnected.
20.8., 14:30 TB, JP Recorded pedestal file 1440073816
20.8., 15:00 MJ, CD Add daquser to the sudoers file on pcbl4sleo4g -
19.8., 15:00 MJ, JP The fixed scintillator of T9 (SC0) is now connected to the control room -
18.8., 14:00 MJ V560 installed, VMEbus A32 base address = 0x800000 -
18.8. Morning TB, CD, MJ, JP Cabling control room, testing detectors Cherenkov volumes down to 0.01 bar using vacuum pumps
17.8., 18:00 MJ, CD, TB To be discussed tomorrow with BI expert:
cable routing
Cherenkov set-up
DWC gas connector
DWC signals (why do we see some in the control room?)
Can we use the scintillators next to the first DWC?
Philippe Carriere will:
* Check the meter of the Cherenkov pump
* Provide us with a connector for the DWC gas supply
* He explained that it is normal to get ~10 Hz signal rate on the DWC
17.8., 18:00 MJ, CD, TB To be discussed tomorrow with Adrian:
What equipment can stay in T9 & counting room during SHIP?
Installation of MNP17 magnet,
how do we control the magnet?
-
17.8, 14:00 TB Installation continued See below how HV (Cable map) and signal are cabled
14.8., all day MJ, TB, CD H/W installed in the T9 zone CD: Timepix not working

phone numbers

Name Phone Function
Philippe Carriere 162545 Beam instrumentation
Adrian Fabich 160345 PS control (if you need beam...)
Michael Jeckel 164710 T9 operation
Lau Gatignon 163554 Everything
Andre Henriques 169774 Safety
Michael Lazzaroni 162407 MNP17 magnet
Jens Spanggaard 163629 gas components?

Cable table

ID Description H/V cable HV PSU port Signal cable TDC channel Scalar channel QDC channel ADC channel
CH0 Upstream (BCA) Che0 Neg-4 91 2-3 3    
CH1 Downstream (BCB) Che1 Neg-5 P-05 2-4 4    
SC0 Fixed beam scintillator - - 4-09 (last meter is patch-cable B-02) 2-0 0    
SC1 9mm Veto scintillator CH0 Neg-6 P-12 2-1 1    
SC2 200x300mm plate scintillator LG10 Neg-7 currently not connected 2-2 2    
DWC0 Fixed beam monitor looking downstream - -
DWC0-U       P-01 1-0      
DWC0-D       P-02 1-1      
DWC0-L       P-03 1-2      
DWC0-R       P-04 1-3      
DWC1 Installed in front of BGO looking upstream LG5 Pos-0
DWC1-U       4-01 1-4      
DWC1-D       4-02 1-5      
DWC1-L       4-03 1-6      
DWC1-R       4-04 1-7      
DWC2 Installed in front of BGO looking upstream LG1 Pos-1
DWC2-U       P-07 1-8      
DWC2-D       P-08 1-9      
DWC2-L       P-09 1-10      
DWC2-R       P-10 1-11      
BGO On DESY table next to Lead Glass LG8 Pos-2 P-11       0
LG0   LG6 Neg-0 4-05     0  
LG8   LG7 Neg-1 4-06     1  
LG3   LG9 Neg-2 4-07     2  
LG4   LG11 Neg-3 4-08     3  
CoDet SC0+SC2+CoCH       2-5 5
CoCH CH0+CH1       2-6 6
Timer 100us pulser         8
Spill CPS spill signal         9

VME Address map

Module Address Address type
v792 0x0 A32
CORBO 0x700000 A24
v560 0x800000 A32
v1290 #1 0x4000000 A32
v1290 #2 0x5000000 A32
v785 0x6000000 A32

-- TimBrooks - 2015-08-17

Edit | Attach | Watch | Print version | History: r38 < r37 < r36 < r35 < r34 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r38 - 2015-09-14 - MarkusJoos
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    BL4S All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback